# Polynomial invariants for $3$-dimensional Leibniz algebras 111 The work is supported by the FCT 2023.08031.CEECIND, UIDB/00212/2020, UIDP/00212/2020, and FAPESP 2023/17918-2

### v1

@@ 4 Automorphisms of $3$-dimensional Leibniz algebras
$\mathcal{L}_{1}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{5}^{2}&\alpha_{5}\alpha_{6}&\alpha_{6}^{2}/2\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
$\mathcal{L}_{2}^{\lambda\neq 0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&0\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1},\alpha_{5}\neq 0;$
$\mathcal{L}_{2}^{0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1},\alpha_{5}\neq 0;$
$\mathcal{L}_{3}$ :::: $g=\left(\begin{array}[]{ccc}1&0&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
$\mathcal{L}_{4}^{\lambda\neq 0}$ :::: $g_{1}=\left(\begin{array}[]{ccc}\frac{\alpha_{6}^{2}+\alpha_{6}\alpha_{9}+\lambda\alpha_{9}^{2}}{\lambda}&\alpha_{2}&\alpha_{3}\\ 0&\frac{\alpha_{6}}{\lambda}+\alpha_{9}&\alpha_{6}\\ 0&-\frac{\alpha_{6}}{\lambda}&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{6}\neq 0,\;\det(g_{1})\neq 0,$
@@
$\mathcal{L}_{4}^{0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{5}(\alpha_{5}+\alpha_{8})&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&0\\ 0&\alpha_{8}&\alpha_{5}+\alpha_{8}\\ \end{array}\right)\text{ for }\alpha_{5}+\alpha_{8}\neq 0;$
$\mathcal{L}_{5}$ :::: $g_{1}=\left(\begin{array}[]{ccc}\alpha_{5}^{2}+\alpha_{6}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&-\alpha_{6}&\alpha_{5}\\ \end{array}\right)\text{ for }a_{5}\neq 0,\;\alpha_{5}^{2}+\alpha_{6}^{2}\neq 0,$
@@
$\mathcal{L}_{6}$ :::: $g_{1}=\left(\begin{array}[]{ccc}\alpha_{1}&\alpha_{4}&0\\ \alpha_{4}&\alpha_{1}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}\neq\alpha_{4}^{2};$
@@
$\mathcal{L}_{7}^{\lambda\neq 0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&\lambda\alpha_{4}&0\\ \alpha_{4}&\alpha_{1}+\alpha_{4}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}+\alpha_{1}\alpha_{4}-\lambda\alpha_{4}^{2}\neq 0;$
$\mathcal{L}_{7}^{0}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&\alpha_{3}\\ \alpha_{4}&\alpha_{1}+\alpha_{4}&-\alpha_{3}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}(\alpha_{1}+\alpha_{4})\neq 0;$
$\mathcal{L}_{8}$ :::: $g=\left(\begin{array}[]{ccc}\alpha_{9}^{2}&0&\alpha_{3}\\ \alpha_{3}\alpha_{9}&\alpha_{9}^{3}&\alpha_{6}\\ 0&0&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{9}\neq 0;$
$\mathcal{L}_{9}$ :::: $g=\left(\begin{array}[]{ccc}1+\alpha_{3}&0&\alpha_{3}\\ \alpha_{3}&1&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{3}\neq-1.$
@@
    Assume $\mathcal{L}=\mathcal{L}_{2}^{\lambda}$ for some $\lambda\in\mathbb{F}$.
    1. (a)
      Let $\lambda\neq 0$. Consequently considering $g_{13}=0$, $g_{23}=0$, $g_{32}=0$ we obtain $\alpha_{7}=0$, $\alpha_{8}=0$, $\alpha_{2}=0$, respectively. Since $\det(g)=\alpha_{1}\alpha_{5}\alpha_{9}\neq 0$, we consequently consider $g_{31}=0$, $g_{33}=0$, $g_{13}=0$ to obtain $\alpha_{4}=0$, $\alpha_{3}=0$, $\alpha_{9}=1$, respectively.
    2. (b)
      Let $\lambda=0$. Since $g_{32}=0$, we have $\alpha_{2}=0$ and $\alpha_{8}=0$. Since $\det(g)=-\alpha_{3}\alpha_{5}\alpha_{7}+\alpha_{1}\alpha_{5}\alpha_{9}$, we have $\alpha_{5}\neq 0$. Since $g_{12}=0$, we have $\alpha_{7}=0$. Thus $\det(g)=\alpha_{1}\alpha_{5}\alpha_{9}\neq 0$ and we consequently consider $g_{13}=0$, $g_{23}=0$ to obtain $\alpha_{4}=0$, $\alpha_{9}=1$, respectively. The required statement is proven.
@@
Given $\mathcal{L}\in{\bf L}_{3}$, the group $\mathop{\rm Aut}(\mathcal{L})$ has dimension $2$, $3$ or $4$ as an affine variety, where each of the possibilities holds for some algebra $\mathcal{L}\in{\bf L}_{3}$.
@@
[…]
@@
$\mathcal{L}_{1}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha,1\},$ $\mathcal{L}_{2}^{\lambda}$ :::: $\delta={\rm diag}\{\alpha,\beta,1\},$
$\mathcal{L}_{3}$ :::: $\delta={\rm diag}\{1,\alpha,1\},$ $\mathcal{L}_{4}^{\lambda}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha,\alpha\},$
$\mathcal{L}_{5}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha,\pm\alpha\},$ $\mathcal{L}_{6}$ :::: $\delta={\rm diag}\{\alpha,\alpha,1\},$
$\mathcal{L}_{7}^{\lambda}$ :::: $\delta={\rm diag}\{\alpha,\alpha,1\},$ $\mathcal{L}_{8}$ :::: $\delta={\rm diag}\{\alpha^{2},\alpha^{3},\alpha\},$
@@ 1. $\bullet$
    $I_{m}(\mathcal{L}_{1})$, $I_{m}(\mathcal{L}_{2}^{\lambda})$, $I_{m}(\mathcal{L}_{7})$ are all equal to $\mathbb{F}[z_{1},\ldots,z_{m}]$,
@@ 3. $\bullet$
    $I_{m}(\mathcal{L}_{4}^{\lambda})=I_{m}(\mathcal{L}_{5})=I_{m}(\mathcal{L}_{8})=\mathbb{F}$,
@@ 1. $\bullet$
    $I_{m}(\mathcal{L}_{1})$, $I_{m}(\mathcal{L}_{2}^{\lambda})$, $I_{m}(\mathcal{L}_{6})$, $I_{m}(\mathcal{L}_{7})$ belong to $\mathbb{F}[z_{1},\ldots,z_{m}]$,
@@ 3. $\bullet$
    $I_{m}(\mathcal{L}_{4}^{\lambda})=I_{m}(\mathcal{L}_{5})=I_{m}(\mathcal{L}_{8})=\mathbb{F}$.
Theorem 10 implies that $z_{1},\ldots,z_{m}\in I_{m}(\mathcal{L})$ for every $\mathcal{L}$ from the following list: $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{3}$, $\mathcal{L}_{7}$, $\mathcal{L}_{9}$. Thus, the statement of the theorem holds for algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{\lambda}$, $\mathcal{L}_{4}^{\lambda}$, $\mathcal{L}_{5}$, $\mathcal{L}_{7}$, $\mathcal{L}_{8}$.
@@
Consider the algebra $\mathcal{L}_{9}$. Using formlas (5.1) we can see that $(x_{r}-y_{r})z_{s}-z_{r}(x_{s}-y_{s})$ lies in $I_{m}(\mathcal{L}_{9})$. For every $\alpha\in\mathbb{F}\backslash\{-1\}$ consider an automorphism
@@
$\begin{array}[]{rcl}\deg_{x^{\prime}}(w)&=&\deg_{x^{\prime}_{1}}(w)+\cdots+\deg_{x_{m}}(w),\\ \deg_{y}(w)&=&\deg_{y_{1}}(w)+\cdots+\deg_{y_{m}}(w),\\ \deg_{z}(w)&=&\deg_{z_{1}}(w)+\cdots+\deg_{z_{m}}(w).\\ \end{array}$
@@ 2. $\bullet$
$x_{1}^{\prime}x_{2}^{\prime}y_{3}<z_{1}y_{2}y_{3},\;\;\;\;x_{1}^{\prime}x_{2}^{\prime}y_{3}<z_{1}x^{\prime}_{2}y_{3},\;\;\text{ and }\;\;y_{1}z_{2}x^{\prime}_{3}<y_{1}x^{\prime}_{2}y_{3}.$
@@
[…]
@@ 6 Trace invariants
$\mathcal{L}_{2}^{\lambda}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=z_{r},$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=(\lambda-1)z_{r},$
@@ 1. $\bullet$
    ${\rm ncl}(\mathcal{L}_{4}^{\lambda})={\rm ncl}(\mathcal{L}_{5})=3$ for all $\lambda\in\mathbb{F}$;

### v2

@@ 3. $\bullet$
$\mathcal{L}_{1}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{5}^{2}&\alpha_{5}\alpha_{6}&\alpha_{6}^{2}/2\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
$\mathcal{L}_{2}^{{\lambda\neq 0}}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&0\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1},\alpha_{5}\neq 0;$
$\mathcal{L}_{3}$ :: $g=\left(\begin{array}[]{ccc}1&0&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{5}\neq 0;$
$\mathcal{L}_{4}^{\lambda\neq 0}$ :: $g_{1}=\left(\begin{array}[]{ccc}\frac{\alpha_{6}^{2}+\alpha_{6}\alpha_{9}+\lambda\alpha_{9}^{2}}{\lambda}&\alpha_{2}&\alpha_{3}\\ 0&\frac{\alpha_{6}}{\lambda}+\alpha_{9}&\alpha_{6}\\ 0&-\frac{\alpha_{6}}{\lambda}&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{6}\neq 0,\;\det(g_{1})\neq 0,$
@@
$\mathcal{L}_{4}^{0}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{5}(\alpha_{5}+\alpha_{8})&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&0\\ 0&\alpha_{8}&\alpha_{5}+\alpha_{8}\\ \end{array}\right)\text{ for }\alpha_{5}+\alpha_{8}\neq 0;$
$\mathcal{L}_{5}$ :: $g_{1}=\left(\begin{array}[]{ccc}\alpha_{5}^{2}+\alpha_{6}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&-\alpha_{6}&\alpha_{5}\\ \end{array}\right)\text{ for }a_{5}\neq 0,\;\alpha_{5}^{2}+\alpha_{6}^{2}\neq 0,$
@@
$\mathcal{L}_{6}$ :: $g_{1}=\left(\begin{array}[]{ccc}\alpha_{1}&\alpha_{4}&0\\ \alpha_{4}&\alpha_{1}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}\neq\alpha_{4}^{2};$
@@
$\mathcal{L}_{7}^{\lambda\neq 0}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{1}&\lambda\alpha_{4}&0\\ \alpha_{4}&\alpha_{1}+\alpha_{4}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}^{2}+\alpha_{1}\alpha_{4}-\lambda\alpha_{4}^{2}\neq 0;$
$\mathcal{L}_{7}^{0}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{1}&0&\alpha_{3}\\ \alpha_{4}&\alpha_{1}+\alpha_{4}&-\alpha_{3}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}(\alpha_{1}+\alpha_{4})\neq 0;$
$\mathcal{L}_{8}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{9}^{2}&0&\alpha_{3}\\ \alpha_{3}\alpha_{9}&\alpha_{9}^{3}&\alpha_{6}\\ 0&0&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{9}\neq 0;$
$\mathcal{L}_{9}$ :: $g=\left(\begin{array}[]{ccc}1+\alpha_{3}&0&\alpha_{3}\\ \alpha_{3}&1&\alpha_{6}\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{3}\neq-1;$
$\mathcal{L}_{10}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{1}&\alpha_{2}&0\\ \alpha_{4}&\alpha_{5}&0\\ 0&0&1\\ \end{array}\right)\text{ for }\alpha_{1}\alpha_{5}-\alpha_{2}\alpha_{4}\neq 0;$
$\mathcal{L}_{11}$ :: $g=\left(\begin{array}[]{ccc}\alpha_{9}^{2}&\alpha_{2}&\alpha_{3}\\ 0&\alpha_{5}&\alpha_{6}\\ 0&0&\alpha_{9}\\ \end{array}\right)\text{ for }\alpha_{5}\alpha_{9}\neq 0.$
@@
    Assume $\mathcal{L}=\mathcal{L}_{2}^{\lambda}$ for some non-zero $\lambda\in\mathbb{F}$. Consequently considering $g_{13}=0$, $g_{23}=0$, $g_{32}=0$ we obtain $\alpha_{7}=0$, $\alpha_{8}=0$, $\alpha_{2}=0$, respectively. Since $\det(g)=\alpha_{1}\alpha_{5}\alpha_{9}\neq 0$, we consequently consider $g_{31}=0$, $g_{33}=0$, $g_{13}=0$ to obtain $\alpha_{4}=0$, $\alpha_{3}=0$, $\alpha_{9}=1$, respectively.
@@
  10. 10.
    Assume $\mathcal{L}=\mathcal{L}_{10}$. Equality $g_{13}=0$ implies that $\alpha_{7}=0$. Applying equality $g_{23}=0$, we can see that $\alpha_{8}=0$. Note that $\alpha_{9}\neq 0$, since $\det(g)\neq 0$. Hence, it follows from $g_{33}=0$ that $\alpha_{3}=\alpha_{6}=0$.
    1. (a)
      Let $\alpha_{9}\neq 1$. Then $g_{13}=0$ implies $\alpha_{1}=\alpha_{4}=0$; hence $\det(g)=0$, a contradiction.
    2. (b)
      Let $\alpha_{9}=1$. Then the required statement holds.
  11. 11.
    Assume $\mathcal{L}=\mathcal{L}_{11}$. Equality $g_{11}=0$ implies that $\alpha_{7}=0$. Applying equality $g_{22}=0$, we can see that $\alpha_{8}=0$. Hence, it follows from equality $g_{33}=0$ that $\alpha_{4}=0$ and $\alpha_{1}=\alpha_{9}^{2}$. The required statement is proven.
@@
Given $\mathcal{L}\in{\bf L}_{3}$, the group $\mathop{\rm Aut}(\mathcal{L})$ has dimension $2\leqslant d\leqslant 5$ as an affine variety, where each of the possibilities holds for some algebra $\mathcal{L}\in{\bf L}_{3}$.
Proof.
For each $\mathcal{L}\in{\bf L}_{3}$, the group $\mathop{\rm Aut}(\mathcal{L})$ is the following affine variety:
$\left\{\Big(b_{1},\ldots,b_{d},f_{1}({\underline{b}}),\ldots,f_{r}({\underline{b}})\Big)\,\Big|\,b_{1},\ldots,b_{d}\in\mathbb{F}\text{ and }h_{1}({\underline{b}})\neq 0,\ldots,h_{s}({\underline{b}})\neq 0\right\}$
for some polynomials $f_{1},\ldots,f_{r},h_{1},\ldots,h_{s}$ from $\mathbb{F}[x_{1},\ldots,x_{d}]$ with $r,s\geqslant 0$, where $\underline{b}$} stands for $(b_{1},\ldots,b_{d})$. Then the dimension of $\mathop{\rm Aut}(\mathcal{L})$ is $d$. Applying this reasoning together with Theorem 10, we complete the proof. ∎
@@
If there are $a_{1},\ldots,a_{n}\in\mathbb{N}$ such that $\delta={\rm diag}\{\alpha^{a_{1}},\ldots,\alpha^{a_{n}}\}$ is an automorphism of $\mathcal{A}$ for all $\alpha\in\mathbb{F}$ and $a_{i_{0}}>0$ for some $1\leqslant i_{0}\leqslant n$, then the algebra of invariants $I_{m}(\mathcal{L})$ lies in $\mathbb{F}[x_{ri}\,|\,1\leqslant r\leqslant m,\;1\leqslant i\leqslant n,\;i\neq i_{0}]$. As an example, if ${\rm diag}\{\alpha,1,\ldots,1\}$ is an automorphism of $\mathcal{A}$ for all $\alpha\in\mathbb{F}$, then $I_{m}(\mathcal{L})$ lies in $\mathbb{F}[x_{ri}\,|\,2\leqslant r\leqslant m,\;1\leqslant i\leqslant n]$.
@@
$\mathcal{L}_{1}$ :: $\delta={\rm diag}\{\alpha^{2},\alpha,1\},$ $\mathcal{L}_{2}^{{\lambda\neq 0}}$ :: $\delta={\rm diag}\{\alpha,\beta,1\},$
$\mathcal{L}_{3}$ :: $\delta={\rm diag}\{1,\alpha,1\},$ $\mathcal{L}_{4}^{\lambda}$ :: $\delta={\rm diag}\{\alpha^{2},\alpha,\alpha\},$
$\mathcal{L}_{5}$ :: $\delta={\rm diag}\{\alpha^{2},\alpha,\pm\alpha\},$ $\mathcal{L}_{6}$ :: $\delta={\rm diag}\{\alpha,\alpha,1\},$
$\mathcal{L}_{7}^{\lambda}$ :: $\delta={\rm diag}\{\alpha,\alpha,1\},$ $\mathcal{L}_{8}$ :: $\delta={\rm diag}\{\alpha^{2},\alpha^{3},\alpha\},$
$\mathcal{L}_{10}$ :: $\delta={\rm diag}\{\alpha,\beta,1\},$ $\mathcal{L}_{11}$ :: $\delta={\rm diag}\{\alpha^{2},\beta,\alpha\},$
@@ 1. $\bullet$
    $I_{m}(\mathcal{L}_{1})$, $I_{m}(\mathcal{L}_{2}^{{\lambda\neq 0}})$, $I_{m}(\mathcal{L}_{7})$, $I_{m}(\mathcal{L}_{10})$ are all equal to $\mathbb{F}[z_{1},\ldots,z_{m}]$,
@@ 3. $\bullet$
    $I_{m}(\mathcal{L}_{4}^{\lambda})=I_{m}(\mathcal{L}_{5})=I_{m}(\mathcal{L}_{8})={I_{m}(\mathcal{L}_{11})}=\mathbb{F}$,
@@ 1. $\bullet$
    $I_{m}(\mathcal{L}_{1})$, $I_{m}(\mathcal{L}_{2}^{{\lambda\neq 0}})$, $I_{m}(\mathcal{L}_{6})$, $I_{m}(\mathcal{L}_{7})$, $I_{m}(\mathcal{L}_{10})$ belong to $\mathbb{F}[z_{1},\ldots,z_{m}]$,
@@ 3. $\bullet$
    $I_{m}(\mathcal{L}_{4}^{\lambda})=I_{m}(\mathcal{L}_{5})=I_{m}(\mathcal{L}_{8})={I_{m}(\mathcal{L}_{11})}=\mathbb{F}$.
Theorem 10 implies that $z_{1},\ldots,z_{m}\in I_{m}(\mathcal{L})$ for every $\mathcal{L}$ from the following list: $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{{\lambda\neq 0}}$, $\mathcal{L}_{3}$, $\mathcal{L}_{7}$, $\mathcal{L}_{9}$, $\mathcal{L}_{10}$. Thus, the statement of the theorem holds for algebras $\mathcal{L}_{1}$, $\mathcal{L}_{2}^{{\lambda\neq 0}}$, $\mathcal{L}_{4}^{\lambda}$, $\mathcal{L}_{5}$, $\mathcal{L}_{7}$, $\mathcal{L}_{8}$, $\mathcal{L}_{10}$, $\mathcal{L}_{11}$.
@@
[…]
@@
For $x_{r}^{\prime}=x_{r}-y_{r}$ we have that $\mathbb{F}[\mathcal{L}^{m}]=\mathbb{F}[x^{\prime}_{r},y_{r},z_{r}\,|\,1\leqslant r\leqslant m]$, and until the end of this proof we consider elements of $\mathbb{F}[\mathcal{L}^{m}]$ as linear combinations of monomials in $x^{\prime}_{r}$, $y_{r}$, $z_{r}$.
@@ 2. $\bullet$
[…]
@@
for all $\alpha\in\mathbb{F}$. Since the coordinate ring of $V$ is $\mathbb{F}[x^{\prime}_{r},z_{r}\,|\,1\leqslant r\leqslant m]$, applying Lemma 8 we conclude the proof of the statement about $\mathcal{L}_{9}$.
@@ 6 Trace invariants
$\mathcal{L}_{2}^{{\lambda\neq 0}}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=z_{r},$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=(\lambda-1)z_{r},$
@@
$\mathcal{L}_{10}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=0,$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=2z_{r},$
$\mathcal{L}_{11}$ $\mathop{\rm tr}(\chi_{r}\chi_{0})=0,$ $\mathop{\rm tr}(\chi_{0}\chi_{r})=0,$
@@ 1. $\bullet$
    ${\rm ncl}(\mathcal{L}_{4}^{\lambda})={\rm ncl}(\mathcal{L}_{5})={{\rm ncl}(\mathcal{L}_{11})}=3$ for all $\lambda\in\mathbb{F}$;